OpenStudy (anonymous):

Solve for x: (x - 9)^2 = 1

OpenStudy (anonymous):

Show work

OpenStudy (anonymous):

x - 9 = pm 1

OpenStudy (anonymous):

FOIL

OpenStudy (anonymous):

\[x-9=1, x-9=-1\] \[x=10,x=8\]

OpenStudy (anonymous):

lmfao

OpenStudy (anonymous):

lol